Physics Shift Report¶
Start Date: 19/09/2017
End Date: 23/09/2017
Start Run: 09618
End Run: 09628
Physics shifter: Chris Rogers
Summary¶
Attempted to ramp magnets; but magnet experts have struggled to ramp. Running straight tracks for alignment and TOF calibration
Elog, CDB¶
Elog is okay
CDB is okay
Online reco¶
Looks okay. Note that tracker has some dead areas; tracker group has been notified. Also note that there are some spills that are coming through onrec as "0"; i.e. run number 0, spill number 0, no recon events. Durga thinks it is "normal" DAQ misfiring type events
Detectors¶
TOF¶
For run 09628
- TOF0 failed calibration 0.09 % out of range 2.21 % Okay 97.7 %
- TOF1 failed calibration 0.0 % out of range 0.0 % Okay 100.0 %
- TOF2 failed calibration 2.73 % out of range 4.91 % Okay 92.36 %
Delta TOF between slabs is attached (RMS is ~twice the TOF resolution). It looks not bad given that there is a calibration coming.
- tof_validation.pdf is the difference between slabs of all formed space points
- tof_validation_raw.pdf is the uncalibrated difference between slabs; fileld is the uncalibrated difference for formed space points; dashed is the uncalibrated difference for all slab hit combinations.
Tracker¶
- 2017-10-21 There is evidence for some dead channels in the tracker. See this email (cut n paste)
- 2017-10-22 A new mapping was developed to try to resolve some of the dead areas and reconstruction proceeded with this new mapping and calibration during the day on 21st October; following this the space point reconstruction in tkd looked to have poor efficiency and few tracks were formed (tkd made tracks in less than 10 % of events that made a good track in upstream and a space point in tof2). It is also noted that at the same time the solenoid field was ramped (so helical tracking) and a new MAUS version went out; but likely cause is the new mapping/calibration. Experts are investigating.
Beamline tuning¶
- There was beamline tuning activity on the night/early morning of 21st/22nd Oct. See these elog posts
Analysis¶
Data rates/etc after cuts. Aim is to get 1e5 "good muons" for each beamline setting, i.e. muons in the "us cut" column.
name | runs | lmc1234 | tof1 | tof2 | time | nevents | us cut | ds cut |
2016-04 1.2 3-140-9659 | [9659] | 1687730 | 675336 | 375824 | 2 hrs 4 mins | 387555 | 45559 | 1683 |
2016-04 1.2 3-140-9660 | [9660] | 1692847 | 687797 | 382610 | 2 hrs 10 mins | 400436 | 47507 | 1796 |
2016-04 1.2 6-140-9662 | [9662] | 57568 | 26791 | 11208 | 0 hrs 5 mins | 15807 | 1846 | 27 |
2016-04 1.2 6-140-9663 | [9663] | 1159737 | 539173 | 224165 | 1 hrs 35 mins | 309935 | 37167 | 1440 |
2016-04 1.2 6-140-9664 | [9664] | 998892 | 747666 | 108394 | 1 hrs 32 mins | 376548 | 18751 | 445 |
2016-04 1.2 6-140-9665 | [9665] | 1375337 | 629355 | 261628 | 1 hrs 52 mins | 353562 | 42292 | 1580 |
2016-04 1.2 6-140-9666 | [9666] | 1041647 | 443838 | 184485 | 1 hrs 2 mins | 227014 | 25857 | 954 |
Predicted Rates¶
Updated by Rogers, Chris almost 6 years ago · 24 revisions